OpenStudy (anonymous):

What is the answer to 7-y-y=-1

OpenStudy (anonymous):

y = 4 ???

OpenStudy (anonymous):

7 - 2y = -1 7 + 1 = 2y 8 = 2* y y = 8/2 y = 4

OpenStudy (anonymous):

@roanna explanation is perfect. I can't explain it better.

OpenStudy (anonymous):

@mtaOS, I hope this isn't a sarcasm XD

OpenStudy (anonymous):

Of course not man. I was serious..... I just can't explain it any better. you were completely right and you wrote each step.

OpenStudy (anonymous):

@mtaOS, thanx! )